Sorry the question wasnt done yet...Find the indicated outputs for
f(x)= 2x2 - 5x
(The second 2 is too the power of)(one of those little 2's above the number)
f(0)=
f(-1)=
f(2)=
HELP PLEASE!!!!!

Answer by stanbon(75887)

f(x)= 2x^2 - 5x
(The second 2 is too the power of)(one of those little 2's above the number)
f(0)= 2(0)^2 - 5*2 = 2*0-5*2 = -10
----------
f(-1)= 2(1)^2 - 5*1 = 2*1 - 5 = -3
--------------
f(2)= 2(2)^2 - 5*2 = 2*4 - 10 = -2
==================
Cheers,
Stan H.